OpenStudy (anonymous):

Choose the equation of the line that contains (4,4) and (-2,0). Y=-2/3x-4/3, y=4/3x+17/3, y=5

OpenStudy (helder_edwin):

the slope of the line that contains both points is \[ m=\frac{4-0}{4-(-2)}=\frac{4}{6}=\frac{2}{3} \]

OpenStudy (anonymous):

thanks

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 5 and y-intercept of -3.

OpenStudy (helder_edwin):

the equation of a line of slope m and y-intercept b is \[ y=mx+b \]

OpenStudy (anonymous):

i kinda get that just cant figure it out

OpenStudy (helder_edwin):

the slope has to be m=5 and the y-intercetp b=-3, so the equation is \[ y=5x+(-3) \] \[ y=5x-3 \]